Please answer in plain text.

  1. Determine the C++ output for the following:

Assume: int num1 = 25,   num2 = 5,   num3 = 10;

        

a.    cout<< “ The product is “ << num1 * num3;

       

        ________________________________________________________

b.     cout<< “ The total of the three numbers is: “ <<(num1+num2+num3);

         ________________________________________________________

c.     cout<< “ The total is: “ <<(num1+254.3373);

       

        _________________________________________________________

d.     cout << num1 << “ +  “    << num2 << “ = ” << num1 + num2;

     

         _________________________________________________________

e.    cout << “ This is how the output should look  | “ << (num1- num3)<< “|”;  

            

_________________________________________________________
Answer #1

a. Output: The product is 250

  • num1*num3 = 25*10 = 250

b. Output: The total of the three numbers is: 40

  • num1+num2+num3 = 25+5+10 = 40

c. Output: The total is: 279.337

  • num1+254.3373 = 25+254.3373 = 279.337

d. Output: 25 + 5 = 30

e. Output:  This is how the output should look  | 15|

  • num1-num2 = 25-10 = 10